# MFAI: A Scalable Bayesian Matrix Factorization Approach to Leveraging Auxiliary Information

### v2

@@ 2.1 The MFAI Model
$\left(\widehat{\boldsymbol{\Theta}},\widehat{\mathbf{F}}\left(\cdot\right)\right)$ $=\underset{\boldsymbol{\Theta},\mathbf{F}\left(\cdot\right)}{\operatorname{arg}\,\operatorname{max}}\;\log{\rm Pr}\left(\mathbf{Y}\mid\boldsymbol{\Theta},\mathbf{F}\left(\cdot\right)\right)$ (8)
@@ 2.2 Fitting the MFAI Model
$\left(\widehat{\boldsymbol{\Theta}},\widehat{F}\left(\cdot\right)\right)$ $=\underset{\boldsymbol{\Theta},F\left(\cdot\right)}{\operatorname{arg}\,\operatorname{max}}\;\log{\rm Pr}\left(\mathbf{Y}\mid\boldsymbol{\Theta},F\left(\cdot\right)\right)$ (12)
@@ 2.2.1 Approximate Bayesian Inference
[…]
@@
Now suppose that we are at the $t$-th step of the variational EM algorithm, and we have obtained the current estimates $\left\{\boldsymbol{\mu}^{(t-1)},{a^{2}}^{(t-1)};\boldsymbol{\nu}^{(t-1)},{b^{2}}^{(t-1)}\right\}$, $\boldsymbol{\Theta}^{(t-1)}=\left\{\tau^{(t-1)},\beta^{(t-1)}\right\}$, and $F^{(t-1)}\left(\cdot\right)$ at the $(t-1)$-th step. To maximize ELBO in the $t$-th E-step, we can update variational parameters as
@@
where $10<s<$ 1 is the so-called shrinkage parameter or learning rate, whose default value is set relatively small ($s=0.1$) in our implementation to avoid potential overfitting. Clearly, information in the auxiliary matrix is gradually incorporated to modulate the prior of $\mathbf{z}$, and the corresponding posterior mean will be updated as $\boldsymbol{\mu}^{(t+1)}={a^{2}}^{(t)}\left(\beta^{(t)}F^{(t)}\left(\mathbf{X}\right)+\tau^{(t)}\mathbf{Y}\boldsymbol{\nu}^{(t)}\right)$ in the $(t+1)$-th step.
@@
1 initialize approximate posterior $q^{(0)}=q\left(\boldsymbol{\mu}^{(0)},{a^{2}}^{(0)};\boldsymbol{\nu}^{(0)},{b^{2}}^{(0)}\right)$ ;
@@ 6       $t\leftarrow t+1$ ;
7
8until convergence criterion satisfied;
@@ 2.2.2 Missing Data
The update for the function $F\left(\cdot\right)$ remains the same as (27) and (28). As for the missing data in auxiliary information, it is clear that only the update steps involving the auxiliary matrix $\mathbf{X}$ need to be reconsidered. Utilizing the rpart package (Therneau and Atkinson, 2022a), any observation with value for the dependent variable (i.e., $\boldsymbol{\mu}_{n}$) and at least one independent variable (i.e., one of $\left\{\mathbf{X}_{n1},\dots,\mathbf{X}_{nC}\right\}$) will participate in the modeling. For each split, the observation with the missing split variable will be split based on the best surrogate variable; if that’s missing, then by the next best, and so on (Therneau and Atkinson, 2022b).
@@ 2.2.4 The Multi-Factor MFAI Model
We now extend to fit the multi-factor MFAI model following Wang and Stephens (2021). To do so, we introduce the variational approximations $\left\{q\left(\mathbf{Z}_{\cdot k}\right),q\left(\mathbf{W}_{\cdot k}\right)\right\}$ for $k=1,\dots,K$, and then optimize ${\rm ELBO}(q(\mathbf{Z}_{\cdot 1},\mathbf{W}_{\cdot 1}),\dots,q(\mathbf{Z}_{\cdot K},\mathbf{W}_{\cdot K});\tau,\beta_{1},\dots,\beta_{K},F_{1}(\cdot),\dots,F_{K}(\cdot))$. Similar to the single-factor case, the optimization can be done by iteratively updating parameters relating to a single factor while keeping others fixed. The updates of a single pair $\left\{\mathbf{Z}_{\cdot k},\mathbf{W}_{\cdot k}\right\}$ are essentially identical to those for fitting the single-factor model aforementioned, except that $\mathbf{Y}$ is replaced with the residuals obtained by removing the estimated effects of the other $K-1$ pairs
@@
1 for $k=1,\dots,K$ do
2       $\boldsymbol{\mu}_{k},a^{2}_{k};\boldsymbol{\nu}_{k},b^{2}_{k};\tau_{k},\beta_{k};F_{k}\left(\cdot\right)\leftarrow$ MFAI_SF$(\mathbf{Y},\mathbf{X})$ ;
@@ 3       $\mathbf{Y}\leftarrow\mathbf{Y}-\boldsymbol{\mu}_{k}\boldsymbol{\nu}_{k}^{\mathrm{T}}$ ;
6return $\widehat{\mathbf{Z}}=\left(\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{K}\right);a_{1}^{2},\dots,a_{K}^{2};\widehat{\mathbf{W}}=\left(\boldsymbol{\nu}_{1},\dots,\boldsymbol{\nu}_{K}\right);b_{1}^{2},\dots,b_{K}^{2}$ ;
@@
1 initialize $\left\{\boldsymbol{\mu}_{1},a^{2}_{1};\boldsymbol{\nu}_{1},b^{2}_{1};\tau_{1},\beta_{1};F_{1}\left(\cdot\right)\right\},\dots,\left\{\boldsymbol{\mu}_{K},a^{2}_{K};\boldsymbol{\nu}_{K},b^{2}_{K};\tau_{1},\beta_{K};F_{K}\left(\cdot\right)\right\}$ using greedy algorithm MFAI_greedy$(\mathbf{Y},\mathbf{X})$ ;
@@
5             $\boldsymbol{\mu}_{k},a^{2}_{k};\boldsymbol{\nu}_{k},b^{2}_{k};\tau_{k},\beta_{k};F_{k}\left(\cdot\right)\leftarrow$ MFAI_SF$(\mathbf{R},\mathbf{X})$ with current estimate of $\left\{\boldsymbol{\mu}_{k},a^{2}_{k};\boldsymbol{\nu}_{k},b^{2}_{k};\tau_{k},\beta_{k};F_{k}\left(\cdot\right)\right\}$ as initialization ;
6
@@
8
9until convergence criterion satisfied;
10return $\widehat{\mathbf{Z}}=\left(\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{K}\right);a_{1}^{2},\dots,a_{K}^{2};\widehat{\mathbf{W}}=\left(\boldsymbol{\nu}_{1},\dots,\boldsymbol{\nu}_{K}\right);b_{1}^{2},\dots,b_{K}^{2}$ ;
@@
A practical issue with matrix factorization is how to select the number of factors $K$. Regularized methods often rely on technology such as cross-validation, which can introduce high computational costs. In contrast, taking advantage of the additive model and the stage-wise manner to fit the $K$-factor model sequentially as shown in Algorithm 2, MFAI can automatically determine $K$ with a little modification to the algorithm. We first set the maximum value of the number of factors $K_{\max}$ and perform the for-loop in the greedy algorithm 2 with $K$ replaced by $K_{\max}$. In this process, if we find the $k$-th factor/loading combination $\boldsymbol{\mu}_{k}\boldsymbol{\nu}_{k}^{\mathrm{T}}$ is very close to zero for one specific $k\in\left\{1,\dots,K_{\max}\right\}$, then we stop the fitting procedure and only use the first $k-1$ factors as the final estimates. This means that the users only need to set $K_{\max}$ sufficiently large, and MFAI can automatically find the suitable $K$ without excessive computational costs. The modified algorithm is summarized in Appendix Section B. It is worth emphasizing that our approach to determining $K$ is very similar to the methods using automatic relevancy determination (ARD) prior (MacKay, 1995; Neal, 1996; Tipping, 1999), such as Babacan et al. (2012), whose key idea is that if the data are consistent with a small absolute value, then the prior precision will be estimated to be large, which results in the shrinkage of the corresponding factor/loading combinations towards zero and hence reduces the rank of the estimate.
@@ 3 Numerical Experiments
Method
R package
Brief description
MFAI
mfair
Variational inference, incorporates auxiliary information via gradient boosting machine
EBMF
flashr
Variational inference for a general empirical Bayes matrix factorization model (Wang and Stephens, 2021)
hardImpute
softImpute
Singular value decomposition (Mazumder et al., 2010)
softImpute
softImpute
Fits a regularized low-rank matrix using a nuclear-norm penalty (Mazumder et al., 2010; Hastie et al., 2015)
CMF
cmfrec
Main matrix and auxiliary matrix share the same latent factors (Singh and Gordon, 2008; Cortes, 2018)
@@ 3.1.1 Imputation Accuracy
The simulation datasets were generated as follows. For all settings, we fixed the number of samples at $N=1,000$, the number of features at $M=1,000$, the number of covariates at $C=3$, and the number of factors at $K=3$. The auxiliary matrix $\mathbf{X}=\left[\mathbf{X}_{\cdot 1},\mathbf{X}_{\cdot 2},\mathbf{X}_{\cdot 3}\right]\in\mathbb{R}^{1,000\times 3}$ was generated from uniform distribution $\mathbf{X}_{nc}\overset{i.i.d.}{\sim}\mathcal{U}\left(-10,10\right)$. Then we generated the means of three latent factors $\mathbf{Z}=\left[\mathbf{Z}_{\cdot 1},\mathbf{Z}_{\cdot 2},\mathbf{Z}_{\cdot 3}\right]\in\mathbb{R}^{1,000\times 3}$ via three functions, $F_{1}\left(\mathbf{x}\right)=\frac{1}{2}x_{1}-x_{2}$, $F_{2}\left(\mathbf{x}\right)=\frac{1}{10}x_{1}^{2}-\frac{1}{10}x_{2}^{2}+\frac{1}{5}x_{1}x_{2}$, and $F_{3}\left(\mathbf{x}\right)=5\sin{\left(\frac{1}{100}x_{3}^{3}\right)}$ ), respectively. We defined the proportion of variance explained (PVE) by $F_{k}(\mathbf{X})$ as ${\rm PVE}_{k}=\frac{\text{Var}\left(F_{k}\left(\mathbf{X}\right)\right)}{\text{Var}\left(F_{k}\left(\mathbf{X}\right)\right)+\beta^{-1}_{k}}$, and controlled ${\rm PVE}_{k}=0.95$ for $k\in\left\{1,2,3\right\}$. Note that $F_{1}\left(\cdot\right)$ is linear, while $F_{2}\left(\cdot\right)$ and $F_{3}\left(\cdot\right)$ are non-linear. We designed $F_{1}\left(\cdot\right)$, $F_{2}\left(\cdot\right)$, and $F_{3}\left(\cdot\right)$ in these forms to examine whether MFAI could flexibly incorporate auxiliary information. The latent loading matrix $\mathbf{W}=\left[\mathbf{W}_{\cdot 1},\mathbf{W}_{\cdot 2},\mathbf{W}_{\cdot 3}\right]\in\mathbb{R}^{1,000\times 3}$ was generated from normal distribution $\mathbf{W}_{mk}\overset{i.i.d.}{\sim}\mathcal{N}\left(0,1\right)$. With $\mathbf{Z}$ and $\mathbf{W}$, we obtained the true value $\mathbf{Y}^{\text{true}}=\mathbf{Z}\mathbf{W}^{\mathrm{T}}$. At last, we added noises to simulate noisy observation $\mathbf{Y}=\mathbf{Y}^{\text{true}}+\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}_{nm}\sim N(0,\tau^{-1})$. Then the PVE by the factors is defined as ${\rm PVE}=\frac{\text{Var}\left(\mathbf{Y}^{\text{true}}\right)}{\text{Var}\left(\mathbf{Y}^{\text{true}}\right)+\tau^{-1}}$. To mimic the real data with a partially observed main matrix, we randomly masked a subset of entries of $\mathbf{Y}$ and denoted their index set as $\Omega^{\text{miss}}$. The remaining entries were considered as observed entries with index set $\Omega^{\text{obs}}$. The missing ratio can then be computed as $\frac{\left|\Omega^{\text{miss}}\right|}{NM}$.
@@ 3.1.3 Computational Efficiency
[…]
Figure 3: Lineplots for the computation timings against data size. In the left panel, we fixed feature size $M$ and varied sample size $N$. In the right panel, we fixed sample size $N$ and varied feature size $M$. Different shapes of the points and colors of the lines represent the different numbers of auxiliary covariates (i.e., $C$) used in the model, respectively.
[…]
@@ 3.2.2 Imputation Accuracy
[…]
@@ 4 Discussion
The auxiliary information is particularly useful to improve matrix factorization when the observed matrix is noisy and sparse. Despite the massive availability of auxiliary information in real-world applications, existing methods largely rely on linear models to connect auxiliary covariates with the main matrix, leading to suboptimal performance. In this article, we propose a scalable Bayesian matrix factorization approach named MFAI to leverage auxiliary information. By integrating the gradient boosted trees with probabilistic matrix factorization, MFAI enables nonlinear modeling of auxiliary covariates and allows the model parameters to be automatically estimated under the empirical Bayes framework, making MFAI adaptive to the complicated connections between the main matrix and auxiliary information. Besides, MFAI naturally inherits several salient features of gradient boosted trees, such as robustness to irrelevant features, immunity to missing values in predictors, and the ability to distinguish useful covariates in the auxiliary information. Under the variational assumption, we developed an efficient algorithm that can simultaneously estimate the number of latent factors and infer the low-rank structure. With our innovations in the model and algorithm designs, our mfair software is effective, stable, and scalable to large datasets. Through comprehensive simulation studies, we showed that MFAI is statistically accurate, especially in the scenario of high sparsity and weak signal strength. We applied MFAI to two real datasets, including one human brain gene expression dataset, and showed that MFAI not only improves the imputation accuracy but also yields biological insights into spatiotemporal gene regulation patterns in the human brain.

### v3

@@ 2.1 The MFAI Model
$=\underset{\boldsymbol{\Theta},\mathbf{F}\left(\cdot\right)}{\operatorname{arg}\,\operatorname{max}}\;\log\int{\rm Pr}\left(\mathbf{Y},\mathbf{Z},\mathbf{W}\mid\boldsymbol{\Theta},\mathbf{F}\left(\cdot\right)\right)\,d\mathbf{Z}\,d\mathbf{W}.$
@@ 2.2 Fitting the MFAI Model
$=\underset{\boldsymbol{\Theta},F\left(\cdot\right)}{\operatorname{arg}\,\operatorname{max}}\;\log\int{\rm Pr}\left(\mathbf{Y},\mathbf{z},\mathbf{w}\mid\boldsymbol{\Theta},F\left(\cdot\right)\right)\,d\mathbf{z}\,d\mathbf{w}.$
@@ 2.2.1 Approximate Bayesian Inference
$\log{\rm Pr}\left(\mathbf{Y}\mid\boldsymbol{\Theta},F\left(\cdot\right)\right)$ $=\log\int{\rm Pr}\left(\mathbf{Y},\mathbf{z},\mathbf{w}\mid\boldsymbol{\Theta},F\left(\cdot\right)\right)\,d\mathbf{z}\,d\mathbf{w}$ (14)
@@
Now suppose that we are at the $t$-th step of the variational EM algorithm, and we have obtained the current estimates $\left\{\boldsymbol{\mu}^{(t-1)},{a^{2}}^{(t-1)};\boldsymbol{\nu}^{(t-1)},{b^{2}}^{(t-1)}\right\}$, $\boldsymbol{\Theta}^{(t-1)}=\left\{\tau^{(t-1)},\beta^{(t-1)}\right\}$, and $F^{(t-1)}\left(\cdot\right)$ at the (t1)(t-1)-th step. To maximize ELBO in the $t$-th E-step, we can update variational parameters as
@@
where $0<s<1$ is the so-called shrinkage parameter or learning rate, whose default value is set relatively small ($s=0.1$) in our implementation to avoid potential overfitting. Clearly, information in the auxiliary matrix is gradually incorporated to modulate the prior of $\mathbf{z}$, and the corresponding posterior mean will be updated as $\boldsymbol{\mu}^{(t+1)}={a^{2}}^{(t)}\left(\beta^{(t)}F^{(t)}\left(\mathbf{X}\right)+\tau^{(t)}\mathbf{Y}\boldsymbol{\nu}^{(t)}\right)$ in the $(t+1)$-th step.
@@
1
21exinitialize approximate posterior $q^{(0)}=q\left(\boldsymbol{\mu}^{(0)},{a^{2}}^{(0)};\boldsymbol{\nu}^{(0)},{b^{2}}^{(0)}\right)$ ;
@@ 7 $t\leftarrow t+1$ ;
8
9until convergence criterion satisfied;
@@ 2.2.2 Missing Data
The update for the function $F\left(\cdot\right)$ remains the same as (27) and (28). As for the missing data in auxiliary information, it is clear that only the update steps involving the auxiliary matrix $\mathbf{X}$ need to be reconsidered. Using the rpart package (Therneau and Atkinson, 2022a), any observation with value for the dependent variable (i.e., $\boldsymbol{\mu}_{n}$) and at least one independent variable (i.e., one of $\left\{\mathbf{X}_{n1},\dots,\mathbf{X}_{nC}\right\}$) will participate in the modeling. For each split, the observation with the missing split variable will be split based on the best surrogate variable; if that’s missing, then by the next best, and so on (Therneau and Atkinson, 2022b).
@@ 2.2.4 The Multi-Factor MFAI Model
We now extend to fit the multi-factor MFAI model following Wang and Stephens (2021). To do so, we introduce the variational approximations $\left\{q\left(\mathbf{Z}_{\cdot k}\right),q\left(\mathbf{W}_{\cdot k}\right)\right\}$ for $k=1,\dots,K$, and then optimize ${\rm ELBO}(q(\mathbf{Z}_{\cdot 1},\mathbf{W}_{\cdot 1}),\dots,q(\mathbf{Z}_{\cdot K},\mathbf{W}_{\cdot K});\tau,\beta_{1},\dots,\beta_{K},F_{1}(\cdot),\dots,F_{K}(\cdot))$. Similar to the single-factor case, the optimization can be done by iteratively updating parameters relating to a single factor while keeping others fixed. The updates of a single pair $\left\{\mathbf{Z}_{\cdot k},\mathbf{W}_{\cdot k}\right\}$ are essentially identical to those for fitting the single-factor model aforementioned, except that $\mathbf{Y}$ is replaced with the residuals obtained by removing the estimated effects of the other K1K-1 pairs
@@
1
21exfor $k=1,\dots,K$ do
[…]
@@ 4 $\mathbf{Y}\leftarrow\mathbf{Y}-\boldsymbol{\mu}_{k}\boldsymbol{\nu}_{k}^{\mathrm{T}}$ ;
7return $\widehat{\mathbf{Z}}=\left(\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{K}\right);a_{1}^{2},\dots,a_{K}^{2};\widehat{\mathbf{W}}=\left(\boldsymbol{\nu}_{1},\dots,\boldsymbol{\nu}_{K}\right);b_{1}^{2},\dots,b_{K}^{2}$ ;
@@
1
[…]
@@
[…]
7
@@
9
10until convergence criterion satisfied;
11return $\widehat{\mathbf{Z}}=\left(\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{K}\right);a_{1}^{2},\dots,a_{K}^{2};\widehat{\mathbf{W}}=\left(\boldsymbol{\nu}_{1},\dots,\boldsymbol{\nu}_{K}\right);b_{1}^{2},\dots,b_{K}^{2}$ ;
@@
A practical issue with matrix factorization is how to select the number of factors $K$. Regularized methods often rely on technology such as cross-validation, which can introduce high computational costs. In contrast, taking advantage of the additive model and the stage-wise manner to fit the $K$-factor model sequentially as shown in Algorithm 2, MFAI can automatically determine $K$ with a little modification to the algorithm. We first set the maximum value of the number of factors $K_{\max}$ and perform the for-loop in the greedy algorithm 2 with $K$ replaced by $K_{\max}$. In this process, if we find the $k$-th factor/loading combination $\boldsymbol{\mu}_{k}\boldsymbol{\nu}_{k}^{\mathrm{T}}$ is very close to zero for one specific $k\in\left\{1,\dots,K_{\max}\right\}$, then we stop the fitting procedure and only use the first k1k-1 factors as the final estimates. This means that the users only need to set $K_{\max}$ sufficiently large, and MFAI can automatically find the suitable $K$ without excessive computational costs. The modified algorithm is summarized in Appendix Section B. It is worth emphasizing that our approach to determining $K$ is very similar to the methods using automatic relevancy determination (ARD) prior (MacKay, 1995; Neal, 1996; Tipping, 1999), such as Babacan et al. (2012), whose key idea is that if the data are consistent with a small absolute value, then the prior precision will be estimated to be large, which results in the shrinkage of the corresponding factor/loading combinations towards zero and hence reduces the rank of the estimate.
@@ 3 Numerical Experiments
Method R package Brief description
MFAI mfair Variational inference, incorporates auxiliary information via gradient boosting machine
EBMF flashr Variational inference for a general empirical Bayes matrix factorization model (Wang and Stephens, 2021)
hardImpute softImpute Singular value decomposition (Mazumder et al., 2010)
softImpute softImpute Fits a regularized low-rank matrix using a nuclear-norm penalty (Mazumder et al., 2010; Hastie et al., 2015)
CMF cmfrec Main matrix and auxiliary matrix share the same latent factors (Singh and Gordon, 2008; Cortes, 2018)
@@ 3.1.1 Imputation Accuracy
[…]
@@ 3.1.3 Computational Efficiency
Finally, we show the computational efficiency of MFAI. We first fixed the sample size $N=5,000$ and varied the number of features $M\in\{1,000,\ 2,000,\ 3,000,\ 4,000,\ 5,000\}$ (the left panel of Figure 3), and then fixed $M=5,000$ and varied $N\in\{1,000,\ 2,000,\ 3,000,\ 4,000,\ 5,000\}$ (the right panel of Figure 3).
Figure 3: Lineplots for the computation timings against data size. In the left panel, we fixed feature size $M$ and varied sample size NN. In the right panel, we fixed sample size $N$ and varied feature size $M$. Different shapes of the points and colors of the lines represent the different numbers of auxiliary covariates (i.e., $C$) used in the model, respectively.
To evaluate the computation time of MFAI, we applied the single-factor MFAI and fixed the number of iteration steps as the same value $20$. Furthermore, the experiments were repeated with different numbers of auxiliary covariates, $C\in\{1,000,\ 2,000,\ 3,000\}$, indicated by different colors in Figure 3.
@@ 3.2.2 Imputation Accuracy
Figure 4: Boxplots comparing the accuracy of different methods in imputing missing entries. These two sets of experiments involve the main matrix $\mathbf{Y}$ that varies from rare information ($\text{training ratio}=0.5$, left) to rich information ($\text{training ratio}=0.9$, right). Accuracy is measured by the difference in each method’s RMSE from the MFAI’s RMSE, then divided by the MFAI’s RMSE, with smaller values indicating higher accuracy. The $y$ axis is plotted on the square-root scale to avoid the plots being dominated by methods performed poorly.
@@ 4 Discussion
The auxiliary information is particularly useful to improve matrix factorization when the observed matrix is noisy and sparse. Despite the massive availability of auxiliary information in real-world applications, existing methods largely rely on linear models to connect auxiliary covariates with the main matrix, leading to suboptimal performance. In this article, we propose a scalable Bayesian matrix factorization approach named MFAI to leverage auxiliary information. By integrating the gradient boosted trees with probabilistic matrix factorization, MFAI enables nonlinear modeling of auxiliary covariates and allows the model parameters to be automatically estimated under the empirical Bayes framework, making MFAI adaptive to the complicated connections between the main matrix and auxiliary information. Besides, MFAI naturally inherits several salient features of gradient boosted trees, such as robustness to irrelevant features, immunity to missing values in predictors, and the ability to distinguish useful covariates in the auxiliary information. Under the variational assumption, we developed an efficient algorithm that can simultaneously estimate the number of latent factors and infer the low-rank structure. With our innovations in the model and algorithm designs, our mfair software is effective, stable, and scalable to large datasets. Through comprehensive simulation studies, we showed that MFAI is statistically more accurate than alternative methods, especially in the scenario of high sparsity and weak signal strength. We applied MFAI to two real datasets, including one human brain gene expression dataset, and showed that MFAI not only improves the imputation accuracy but also yields biological insights into spatiotemporal gene regulation patterns in the human brain.